# Automated Model Evaluation for Object Detection via Prediction Consistency and Reliability

### v1

@@ B.2 Hyperparameter Analysis
We conduct experiments with varying hyperparameters for both tuning and sensitivity analysis. Unless otherwise specified, all experiments use RetinaNet [31] with a ResNet-50 [17] backbone evaluated on the corruption-based meta-dataset. Performance is averaged over vehicle and pedestrian detection; for example, $11.86=(13.50+10.23)/2$ and $4.99=(6.43+3.56)/2$ correspond to the performance of BoS [45] and our proposed PCR, respectively.
Confidence Threshold is a reference value used to assess whether a given prediction is considered confident, employed both in the consistency and reliability scores. Figure B.2(a) illustrates the effect of varying the confidence threshold on mAP estimation. Based on these results, we set $c=0.5$ as the confidence threshold throughout the experiments.
Floor Value is used in the reliability score, which provides a lower bound to low-confidence predictions, ensuring a minimal contribution from all predictions. Figure B.2(b) illustrates the effect of varying the floor value on mAP estimation, averaged over vehicle and pedestrian detection. Based on these results, we set $\alpha=0.2$ as the floor value throughout the experiments.
Scale Parameters control the sensitivity of the sigmoid function, which modulates the contribution of confidence to the consistency and reliability scores. For consistency, the scale parameter $k_{C}$ is negative because the consistency score is negatively correlated with mAP, while for reliability, the scale parameter $k_{R}$ is positive because the reliability score is positively correlated with mAP. Figure B.2(c–d) illustrates the effect of varying the scale parameters on mAP estimation, averaged over vehicle and pedestrian detection. Based on these results, we set $k_{C}=-60$ and $k_{R}=10$ as the scale parameters throughout the experiments.

### v2

@@ B.2 Hyperparameter Analysis
We conduct experiments with varying hyperparameters for both tuning and sensitivity analysis. Unless otherwise specified, all experiments use RetinaNet [31] with a ResNet-50 [17] backbone evaluated on the corruption-based meta-dataset, with each analysis performed using the scores to which the hyperparameter applies. Performance is averaged over vehicle and pedestrian detection; for example, from Tables 1 and 2, $11.86=(13.50+10.23)/2$ and $4.77=(6.57+2.96)/2$ correspond to the performance of BoS [45] and our proposed PCR, respectively.
Confidence Threshold is a reference value used to assess whether a given prediction is considered confident, employed both in the consistency and reliability scores. Figure B.2(a) illustrates the effect of varying the confidence threshold on mAP estimation. Based on these results, we set $c=0.4$ as the confidence threshold throughout the experiments.
Floor Value is used in the reliability score, which provides a lower bound to low-confidence predictions, ensuring a minimal contribution from all predictions. Figure B.2(b) illustrates the effect of varying the floor value on mAP estimation, averaged over vehicle and pedestrian detection. Based on these results, we set $\alpha=0.1$ as the floor value throughout the experiments.
Scale Parameters control the sensitivity of the sigmoid function, which modulates the contribution of confidence to the consistency and reliability scores. For consistency, the scale parameter $k_{C}$ is negative because the consistency score is negatively correlated with mAP, while for reliability, the scale parameter $k_{R}$ is positive because the reliability score is positively correlated with mAP. Figure B.2(c–d) illustrates the effect of varying the scale parameters on mAP estimation, averaged over vehicle and pedestrian detection. Based on these results, we set $k_{C}=-10$ and $k_{R}=20$ as the scale parameters throughout the experiments.